Assume a demand equation: Qd = 9-0.1p-Po + 0.01 pcrm + 0.000!Y; and a supply equation: as 0.1p - 0.02Pfg + 0.01N + 0.01T-0.1w where: p price of lobster Pb price of butter $3 pcm= price of crab meat-S200 Y consumer income $20,000 Q- quantity in thousands of units Pigprice of fishing gear $450 N number of firms 700 T- index of technology 300 w wage rate $10 The equilibrium price of lobster is (enter your response as a whole number).

The equilibrium price of lobster is ​$50.

What is the equilibrium quantity is ______ thousand units ​(enter your response rounded to one decimal place​).
Answer #1

Demand equation

Qd = 9-0.1p-pb+0.01pcm+0.0001Y

Substitute the given values in the demand equation

Qd = 9 - 0.1p -3+ 0.01(200) + 0.0001(20,000)

Qd = 9-0.1p-3+2+2

Qd = 10-0.1p

Supply equation

Qs = 0.1p-0.02pfg+0.01N+0.01T-0.1w

Substitute the given values in the supply equation

Qs = 0.1p-0.02(450) +0.01(700) +0.01(300)-0.1(10)

Qs = 0.1p-9+7+3-1

Qs = 0.1p

At market equilibrium, Demand = Supply

Qd = Qs

10-0.1p = 0.1p

10 = 0.1p+0.1p

10=0.2p

P= 10/0.2

P=50

The equilibrium price of lobster is $50

Substitute P = $50 in Qs

Qs = 0.1p

Q = 0.1(50)

Q = 5

Equilibrium quantity = 5